# Can ChatGPT Enable ITS? The Case of Mixed Traffic Control via Reinforcement Learning

### v1

@@ Abstract
The surge in Reinforcement Learning (RL) applications in Intelligent Transportation Systems (ITS) has contributed to its growth as well as highlighted key challenges. However, defining objectives of RL agents in traffic control and management tasks, as well as aligning policies with these goals through an effective formulation of Markov Decision Process (MDP), can be challenging and often require domain experts in both RL and ITS. Recent advancements in Large Language Models (LLMs) such as GPT-4 highlight their broad general knowledge, reasoning capabilities, and commonsense priors across various domains. In this work, we conduct a large-scale user study involving 70 participants to investigate whether novices can leverage ChatGPT to solve complex mixed traffic control problems. Three environments are tested, including ring road, bottleneck, and intersection. We find ChatGPT has mixed results. For intersection and bottleneck, ChatGPT increases number of successful policies by 150% and 136% compared to solely beginner capabilities, with some of them even outperforming experts. However, ChatGPT does not provide consistent improvements across all scenarios.
@@ I INTRODUCTION
Large Language Models (LLMs) represent a significant advancement in artificial intelligence. Their usefulness as a general-purpose tool is anticipated to have a profound societal impact. Experiments with LLMs such as GPT-4 [1, 2], LLaMA [3], and PALM2 [4] demonstrate their strong reasoning and common sense abilities across domains such as math, science, and coding. To achieve this success, LLMs leverage Reinforcement Learning from Human Feedback (RLHF) [5, 6] to solve the alignment problem, i.e., follow user intent by fine-tuning them on human feedback. While there is a strong focus on improving LLMs though Reinforcement Learning (RL), the possibility of leveraging LLMs to assist RL problems is in nascent stages.
@@
Over the last decade, RL has been adopted to address complex control problems in ITS such as traffic management and autonomous driving [8, 9]. As autonomous agents, including vehicles and traffic lights, become more prevalent [10] in ITS, they introduce new challenges and opportunities. One emerging topic is mixed traffic control that uses RL-empowered robot vehicles (RVs) to regulate human-driven vehicles (HVs), thus improving the overall traffic flow [11, 12, 13, 14]. This surge in research interest has attracted a broader audience to participate in the topic, resulting in a growing demand for creative decision-making and control strategies enabled by RL. However, the initial technical barrier, specifically formulating MDP components to align with a control strategy, poses a challenge. LLMs, with their broad knowledge, commonsense priors, and creative capacity, show promise in reducing these barriers and simplifying the process.
@@ II Related Work
Recent works have studied how Large Language Models (LLMs) can be leveraged in a variety of RL tasks. In robotics, LLMs are utilized in planning and navigation by representing the robotic agent’s input (state) in natural language, additionally incorporating the input with visual or raw sensor data for grounding. This approach demonstrates high data efficiency and generalization to unseen environments [16, 17].
@@ III-B2 Bottleneck
$\omega_{\text{precise}}=\left\{\overline{X}_{HV},\overline{V}_{HV},\overline{X}_{RV},\overline{V}_{RV},o_{20}\right\}\text{,}$ (3)
@@ IV-A Participants and Cohorts
Figure 2: Example state space from novice (left) and reward function from novice with ChatGPT (right) for ring. The reward function contains complex terms generated by ChatGPT. ChatGPT provides sound reasoning for using the reward function metrics (see Sec. V-A for details).
We recruit 70 graduate students as participants ($54$ male, $16$ female). The male/female participants have a median age of $23$/$22$ and average $4.93$/$4.26$ years of driving experience.
@@ IV-B ChatGPT Setup
The study group participants use GPT-4 with $8$k context length and temperature $=0.7$, a static model provided by OpenAI with no additional fine-tuning. To simulate the practical use of ChatGPT and enable in-context learning, a chat interface is employed using TypingMind [26].
@@ IV-C Manuscript
We develop two versions of manuscripts [^1] https://github.com/tmvllrrl/its-study (for different cohorts) with the following sections. Study duration is $95$ minutes.
@@ IV-C1 General Instructions and Background Questions
The general instructions outline participant conduct for the study’s duration. Participants complete the background questions discussed in Sec. IV-A.
@@ IV-C4 Mixed Traffic Environment Descriptions and Questions
Each of the three traffic environments (shown in Fig. 1; details in Sec. III-B) receives a general description, problem explanation, and the mixed traffic objective of the RVs. The general description includes details such as the number of RVs present, general flow behavior of traffic, ratio of RVs to human-driven vehicles. We provide a supplementary video [^2] https://youtu.be/qTqgfl76FAo demonstrating the mixed traffic environments. We ask the participants three questions per environment: the state space, the reward function, and briefly explain the rationale behind the reward function. The state spaces participants provide are observation spaces; however, we solely use “state space” to prevent additional complexity/avoid confusion.
@@ V RESULTS
Figure 3: Normalized reward curves during training are shown for the three traffic control environments. Each environment consists of three curves for the control group (Novices), the study group (Novices + ChatGPT), and an expert. For Novices and Novices + ChatGPT, the solid line indicates the average for all participants, with the shaded region representing variance. In both groups across all networks, average rewards increase over the course of training. This validates that both groups are able to develop state spaces and reward functions that are trainable using RL.
@@ V-A Participant Answers Analysis
We find that ChatGPT [25, 1] impacts provided state spaces/reward functions by significantly decreasing invalid answers over using prior knowledge. We consider an answer valid if it contains metrics provided in the metrics bank or have well-defined explanations. Table II presents the number of valid/invalid answers for all three environments. On average, only $61$% of answers from the control group are valid, while $82$% are valid for the study group, a $21$% increase. This illustrates ChatGPT’s capabilities in guiding participants to create valid state spaces/reward functions.
@@ VI CONCLUSIONS
In the future, we aim to advance this study in several directions. First, we are interested in testing whether ChatGPT can assist the design of other ITS applications such as traffic state estimation and simulation [28, 29, 30, 31, 32, 33, 34], vehicle motion and control [35, 36, 37, 38], and robust driving [39, 40, 41]. Second, we are also interested in studying mixed traffic control under other mobility topics such as traffic safety [42] and micromobility [43]. Finally, we would like to test our approach on other RL tasks, including graph sparsification [44] and quality species selection [45].

### v2

@@ Abstract
The surge in Reinforcement Learning (RL) applications in Intelligent Transportation Systems (ITS) has contributed to its growth as well as highlighted key challenges. However, defining objectives of RL agents in traffic control and management tasks, as well as aligning policies with these goals through an effective formulation of Markov Decision Process (MDP), can be challenging and often require domain experts in both RL and ITS. Recent advancements in Large Language Models (LLMs) such as GPT-4 highlight their broad, general knowledge, reasoning capabilities, and commonsense priors across various domains. In this work, we conduct a large-scale user study involving 70 participants to investigate whether novices can leverage ChatGPT to solve complex mixed traffic control problems. The participants’ task is to develop the state space and reward function for three RL mixed traffic control environments, including ring road, bottleneck, and intersection. We find ChatGPT has mixed results. For intersection and bottleneck, ChatGPT increases number of successful policies by 150% and 136% compared to solely beginner capabilities, with some of them even outperforming experts. However, ChatGPT does not provide consistent improvements across all scenarios.
@@ I INTRODUCTION
Large Language Models (LLMs) represent a significant advancement in artificial intelligence. Their usefulness as a general-purpose tool is anticipated to have a profound societal impact. Experiments with LLMs such as GPT-4 [1, 2], LLaMA [3], and PALM2 [4] demonstrate their strong reasoning and common sense abilities across domains such as math, science, and coding [1, 2, 3, 4, 5, 6]. To achieve this success, LLMs leverage Reinforcement Learning from Human Feedback (RLHF) [7, 8] to solve the alignment problem, i.e., follow user intent by fine-tuning them on human feedback. While there is a strong focus on improving LLMs through Reinforcement Learning (RL), leveraging LLMs to assist RL problems is in nascent stages.
@@
Over the last decade, RL has been adopted to address complex control problems in ITS such as traffic management and autonomous driving [10, 11]. As autonomous agents, including vehicles and traffic lights, become more prevalent [12] in ITS, they introduce new challenges and opportunities. One emerging topic is mixed traffic control that uses RL-empowered robot vehicles (RVs) to regulate human-driven vehicles (HVs), thus improving the overall traffic flow  [13, 14, 15]. This surge in research interest has attracted a broader audience to participate in the topic, resulting in a growing demand for creative decision-making and control strategies enabled by RL. However, the initial technical barrier, specifically formulating MDP components to align with a control strategy, poses a challenge. LLMs, with their broad knowledge, commonsense priors, and creative capacity, show promise in reducing these barriers and simplifying the process.
@@ II Related Work
Recent works have studied how Large Language Models (LLMs) can be leveraged in a variety of RL tasks. In robotics, LLMs are utilized in planning and navigation by representing the robotic agent’s input (state) in natural language, additionally incorporating the input with visual or raw sensor data for grounding. This approach demonstrates high data efficiency and generalization to unseen environments [17, 18]. Similarly, using LLMs to guide exploration during training can also help achieve greater sample efficiency [19].
@@ III-B2 Bottleneck
$\text{s}=\left\{\overline{X}_{HV},\overline{V}_{HV},\overline{X}_{RV},\overline{V}_{RV},o_{20}\right\}\text{,}$ (3)
@@ IV-A Participants and Cohorts
Figure 2: Example prompting by two novices, Novice A and Novice B, with ChatGPT response excerpts for the ring environment. The novices’ prompts are given exactly. Only conversation portions are provided for brevity purposes. Novice A, unlike Novice B and most other participants, starts the conversation by asking if a certain state space is good to which ChatGPT responds the state space might be ”insufficient.” Novice B follows more closely participants’ trend to completely rely on ChatGPT for the state space after providing context of the ring environment. Both novices heavily rely on ChatGPT for the reward function; however, Novice B prompts ChatGPT to include time-to-collision based on prior conversation with ChatGPT.
Figure 3: Example state space from novice (left) and reward function from novice with ChatGPT (right) for ring. The novice without ChatGPT’s reward function heavily uses provided metrics, while using ChatGPT allows the reward function to contain complex, generated terms. ChatGPT provides sound reasoning for using the reward function metrics (see Sec. V-A for details).
We recruit 70 graduate students as participants ($54$ male, $16$ female). The male/female participants have a median age of $23$/$22$ and average $4.93$/$4.26$ years of driving experience. Participants come from a machine learning or artificial intelligence course with minimal coverage of reinforcement learning (RL) taught by the same professor. At the study’s date, artificial intelligence participants have slightly more RL knowledge; however, we find no discernible difference between the two courses’ results. We attribute this to the manuscript and ChatGPT’s assistance.
Control Study Control Study Control Study
Valid 25 28 24 27 21 24
Invalid 13 4 14 5 17 8
Total 38 32 38 32 38 32
@@ IV-B ChatGPT Setup
The study group participants use GPT-4 with $8$k context length and temperature $=0.7$, a static model with no additional fine-tuning. The March 14th, 2023 version of GPT-4 is used. To simulate the practical use of ChatGPT, a fresh chat interface is provided to participants using TypingMind [28]. The participants use the same conversation with ChatGPT for the study to enable in-context learning.
@@ IV-C Manuscript
We develop two versions of manuscripts [^1] https://github.com/tmvllrrl/its-study (for different cohorts) with the following four sections. Study duration is $95$ minutes. Within the $95$ minutes, participants are required to spend the first $25$ minutes reading the first three sections outlined below, and then given the remaining time to look over the entire manuscript. This is to ensure a higher probability of participants reading over the provided material, potentially leading to better quality state spaces and reward functions for the mixed traffic environments.
@@ IV-C1 General Instructions and Background Questions
The general instructions outline participant conduct for the study’s duration. and warn participants that rule violations comprise their answers’ integrity and result in their dismissal. Participants complete the background questions discussed in Sec. IV-A.
@@ IV-C4 Mixed Traffic Environment Descriptions and Questions
Each of the three traffic environments (shown in Fig. 1; details in Sec. III-B) receives a general description, problem explanation, and the mixed traffic objective of the RVs. The general description includes details such as the number of RVs present, general flow behavior of traffic, ratio of RVs to human-driven vehicles. We provide a supplementary video [^2] https://youtu.be/qTqgfl76FAo demonstrating the mixed traffic environments. We ask the participants three questions per environment: the state space, the reward function, and briefly explain the rationale behind the reward function. We ask participants to explain their reward function rationale to encourage deeper understanding from the participants. The state spaces participants provide are observation spaces; however, we solely use “state space” to prevent additional complexity/avoid confusion.
@@ V-A Participant Answers Analysis
TABLE III: Number of valid and invalid answers for each traffic environment for the control and study groups.
We find that ChatGPT [27, 1] impacts provided state spaces/reward functions by significantly decreasing invalid answers over using prior knowledge. We consider an answer valid if it contains metrics provided in the metrics bank or have well-defined explanations. Table III presents the number of valid/invalid answers for all three environments. On average, only $61$% of answers from the control group are valid, while $82$% are valid for the study group, a $21$% increase. This illustrates ChatGPT’s capabilities in guiding participants to create valid state spaces/reward functions.
@@
We provide example prompting by two novices (labeled Novice A and Novice B) from the study group for the ring environment’s state space and reward function in Fig. 2. Novice A deviates from the general trend of participants by asking ChatGPT if a provided state space is good for the ring environment (to which ChatGPT correctly assumes the state space may be insufficient). The general trend of participants is showcased with Novice B where Novice B provides context for the ring environment, then asks what a state space might be for the environment. Additionally, both novices more closely follow participants’ general trend of heavily relying on ChatGPT by asking ChatGPT for a good reward function. Novice B demonstrates a common behavior with their reward function prompting by wanting to have a specific metric (for Novice B, time-to-collision) in the function.
@@
Figure 4: Average normalized reward curves during training are shown for the three traffic control environments. Each environment consists of three curves for the control group (Novices), the study group (Novices + ChatGPT), and an expert. For Novices and Novices + ChatGPT, the solid line indicates the average for all participants, with the shaded region representing variance. In both groups across all networks, average rewards increase over the course of training. This validates that both groups are able to develop state spaces and reward functions that are trainable using RL.
@@ VI CONCLUSIONS
In the future, we intend to expand the study to include more participants, traffic control environments, and Large Language Models (LLMs). A comparative analysis of various LLMs could also be an interesting direction. Further, fine-tuning existing LLMs to be able to serve the ITS community by reducing the human effort required is another avenue that we want to explore.
Acknowledgement
This research is supported by NSF IIS-2153426. The authors would like to thank NVIDIA and the University of Tennessee, Knoxville for their support. The authors would also like to thank Heath Mitchell for GPT-4 API access.